OpenStudy (ammarah):

Use the sum/difference identities to solve using the given information. Given sinx=rad 2/5, where pi/2

OpenStudy (ammarah):

@tkhunny

OpenStudy (tkhunny):

Have you considered stating the identity? \(\sin(x+y) = \sin(x)\cos(y)+\cos(x)\sin(y)\)

OpenStudy (ammarah):

yes i know it

OpenStudy (tkhunny):

Use the Pythagorean Theorem and the Quadrant Hints to find the cosine of x.

OpenStudy (ammarah):

ok

OpenStudy (tkhunny):

What do you get?

OpenStudy (ammarah):

1 sec

OpenStudy (tkhunny):

Do the same for the y in Quadrant III

OpenStudy (tkhunny):

Gaaa!!! You'll never make the State Finals like that!

OpenStudy (ammarah):

? for cos x is it in quadrant 1?

OpenStudy (tkhunny):

Why would it be? We are told explicitly that x in in II.

OpenStudy (ammarah):

do i use sin x?

OpenStudy (ammarah):

i got rad 23 for

OpenStudy (ammarah):



OpenStudy (tkhunny):

We're trying that again. 

OpenStudy (tkhunny):

Very good. What is the cosine of that angle?

OpenStudy (ammarah):

rad 23 /5